OpenStudy (anonymous):

x + 5y = -2 2x + y = 5 The point of intersection of the lines has a y-coordinate of _____. -1 1 3

OpenStudy (mathmale):

Choose your method: elimination by addition/subtraction, substitution, graphing.

OpenStudy (anonymous):

I think it is 3 because if you add the two equations using the Addition method. x + 5y = -2 2x + y = 5 ---------- 3x + 6y = 3 IS this correct?

OpenStudy (mathmale):

Your goal is to eliminate one of the variables. Unfortunately, what you have done has not eliminated either x or y. Think this through.

OpenStudy (anonymous):

Okay, So then x + 5y = -2 2x + y = 5 ---------- -1x + 4y = -7 Would this be correct?

OpenStudy (mathmale):

Have you eliminated either x or y?

OpenStudy (mathmale):

Hint: Multiply every term of the first equation by -2. Results?

OpenStudy (mathmale):

-2(x + 5y = -2) = ?

OpenStudy (anonymous):

From what I know, the equation in the Parentheses is solved with the solution being -2, Though the variables have not yet been solved for it. Based on what I know, would the equation be simplified to -2 x -2?

rebeccaxhawaii (rebeccaxhawaii):

still need help?

OpenStudy (mathmale):

Rather than respond to y our comment, I ask you again to multiply the first equation by -2 and to ask you to think about why this may be a good approach. x + 5y = -2 2x + y = 5 -2(x+5y=-2) = ? Add your result to the second equation and see what happens to the variable x.